Above what Fe2+ concentration will Fe(OH)2 precipitate from a buffer solution that has a pH of 8.14 ? The ?sp of Fe(OH)2 is 4.87×10−17.
Answer #1

PH = 0.14 pol = 14 -pH = 5.86 [0H) = 1.38 x 106m Fe(OH), - fet + 204- 5 25 [Fe2+] = S [or] =25 + 1.38 x 10 y 128x100 Ksp = [